# Unlocking Chemical Insights: Superior Molecular Representations from Intermediate Encoder Layers

### v1

@@ Appendix C Layer-wise Performance
Figure 58: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 59: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Appendix G OrbV3-Conservative: Fixed Embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the OrbV3-Conservative model. Each plot illustrates the relationship between fixed embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 60: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 61: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 62: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 63: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 64: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 65: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 66: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 67: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 68: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 69: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 70: Fixed embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
Appendix H Pos-EGNN: Fixed Embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the Pos-EGNN model. Each plot illustrates the relationship between fixed embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 71: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 72: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 73: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 74: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 75: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 76: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 77: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 78: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 79: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 80: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 81: Fixed embeddings vs. finetuned performance for every layer of Pos-EGNN.

### v2

@@ Appendix C Layer-wise Performance
Figure 59: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Figure 60: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Direct.
@@
Appendix G OrbV3-Conservative: Frozen embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the OrbV3-Conservative model. Each plot illustrates the relationship between frozen embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 61: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 62: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 63: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 64: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 65: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 66: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 67: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 68: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 69: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 70: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
@@
Figure 71: Frozen embeddings vs. finetuned performance for every layer of OrbV3-Conservative.
Appendix H Pos-EGNN: Frozen embedding vs. Finetuned Scatter Plots
@@
This appendix section presents a series of scatter plots dedicated to the Pos-EGNN model. Each plot illustrates the relationship between frozen embedding performance and full finetuning performance across the model’s different layers for a specific downstream task. The plots were ordered by Pearson correlation.
@@
Figure 72: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 73: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 74: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 75: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 76: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 77: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 78: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 79: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 80: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 81: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.
@@
Figure 82: Frozen embeddings vs. finetuned performance for every layer of Pos-EGNN.